OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (2, -1). Which equation defines circle P? (x + 2)2 + (y – 1)2 = 64 (x – 2)2 + (y + 1)2 = 64 (x – 2)2 + (y + 1)2 = 8 (x + 2)2 + (y – 1)2 = 8

OpenStudy (anonymous):

please helpppppp!!!!! I'm soooo stuck on how to solve and find equations!

OpenStudy (anonymous):

Use the fact that circles are loci of points equidistant from their center, and Pythagoras' theorem to calculate said distance.

OpenStudy (anonymous):

so you mean I would use this theorem? a^2 + b^2= c^2

OpenStudy (anonymous):

I don't really undestand how, can you please lead me through it @Cheese3.1416

jimthompson5910 (jim_thompson5910):

The general equation of any circle is (x-h)^2 + (y-k)^2 = r^2 where (h,k) is the center and r is the radius

jimthompson5910 (jim_thompson5910):

In your case, the given center is (2, -1) so (h,k) = (2, -1) which means h = 2 and k = -1

jimthompson5910 (jim_thompson5910):

You're also given a radius of 8, so r = 8

jimthompson5910 (jim_thompson5910):

(x-h)^2 + (y-k)^2 = r^2 (x-2)^2 + (y-(-1))^2 = r^2 ... plug in h = 2 and k = -1 (x-2)^2 + (y+1)^2 = r^2 ... simplify (x-2)^2 + (y+1)^2 = 8^2 ... plug in r = 8 (x-2)^2 + (y+1)^2 = 64 ... square 8 to get 64

OpenStudy (anonymous):

ohhhhh, ok that's where the pythagoreans theorem comes into play; that's the basis correct? Then you input the information in said places... I see.

jimthompson5910 (jim_thompson5910):

So the equation of the circle with center (2, -1) with a radius of 8 is (x-2)^2 + (y+1)^2 = 64

jimthompson5910 (jim_thompson5910):

yes basically you can think of any point on the circle as a point on a right triangle like this

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

since a^2 + b^2 = c^2 by the pythagorean theorem, we can say this  which leads us to x^2 + y^2 = r^2